# Down with the Hierarchy: The ‘H’ in HNSW Stands for “Hubs”

### v2

@@ 2. Background: Similarity Search & HNSW
2.1. HNSW Overview
With this formalization of the ANNS problem, we will now briefly review the key elements of the HNSW algorithm, which is the central focus of our benchmarking study. As we alluded to previously, HNSW builds off of prior work in navigable small world graph indexes introduced in (Malkov et al., 2014). Small world graphs are a well-studied phenomenon in both computing and the social sciences and are primarily defined by the fact that the average length of a shortest path between two vertices is small (typically scaling logarithmically with the number of nodes in the network) (Travers and Milgram, 1977; Watts and Strogatz, 1998; Kleinberg, 2000). Small world graphs are also often characterized by the presence of well-connected hub nodes which we discuss further in the next section.
@@ 2.1. HNSW Overview
32:     Return $M$ nodes based on selection heuristic in  (Arya and Mount, 1993)
@@ 2.2. HNSW Search Algorithm
Specifically, the HNSW index is constructed in an iterative fashion. For a newly inserted element $x$, the algorithm will randomly select a maximum layer $l$ and then insert the new point into every layer up to $l$. This randomized process is executed with an exponentially decaying probability distribution such that, in expectation, each subsequent layer has exponentially more nodes than its predecessor. Within a layer, HNSW greedily adds edges between $x$ and its $M$ closest neighbors (where $M$ is a hyperparameter) where the neighbors consist of previously inserted points. This process then repeats in the subsequent layer below using the closest neighbors found in the prior graph as entry points. Through this process, the top layer of the hierarchy will be the coarsest directed graph, consisting of the fewest nodes and edges, and the bottom layer will be the densest and contain all of the nodes, each with connections to (up to) $M$ neighbors. As an additional, and important, optimization, HNSW also implements the pruning heuristic of (Arya and Mount, 1993) that will prune an edge from $u$ to $v$ if there exists another edge from $u$ to a neighbor $w$ of $v$ such that the distance from $u$ to $w$ is less than that of $u$ to $v$.
Figure 2. Visualization of the HNSW search index. We benchmark the full hierarchical graph search procedure against simply searching in the bottom base layer.
3. Related Work
3.1. Near-Neighbor Benchmarks
Due to the importance of the $k$-NNS problem, there have been several large-scale near-neighbor benchmarks in recent years. The original ANN Benchmarks  (Aumüller et al., 2018) made significant impact because it provided the first standard evaluation of near-neigh3bor search algorithms. Since its inception, the benchmark has grown in scope from a handful of algorithms to include over 30 methods on 9 datasets. However, most of the datasets included in ANN Benchmarks are relatively small by modern standards at around one million points. To reflect the growing need for high-dimensional, large-scale embedding search, NeurIPS hosted a competition track known as Big ANN Benchmarks (Simhadri et al., 2022), which evaluates algorithms on five billion-scale datasets. Graph-based indexes easily and consistently place among the top three algorithms in each benchmark. However, these benchmarks are not sufficient for our purposes because they focus exclusively on overall throughput metrics and do not report tail latency numbers such as the 99th percentile, which is important in our case for teasing apart the impact of hierarchy in HNSW.
Hierarchy Studies: Hierarchical structures are considered a fundamental component of graph-based near-neighbor algorithms. As a result, numerous high-performing algorithms use some form of hierarchy including HNSW (Malkov and Yashunin, 2016), ONNG (Iwasaki and Miyazaki, 2018) PANNG (Iwasaki, 2016), and HCNNG (Munoz et al., 2019). However, this practice has recently come under scrutiny. Dobson et al. (2023) show that the hierarchy may be unnecessary for some workloads, based on the observation that HNSW under-performs both HCNNG (which uses a shallower hierarchy) and DiskANN (Jayaram Subramanya et al., 2019) (which does not use a hierarchy). Lin and Zhao (2019) present results on synthetic data showing that the hierarchy is only beneficial for low-dimensional data ($d<32$), but analyze only a limited number of small real-world datasets and – most importantly – do not explain why the hierarchy fails to provide value. In this work, we aim to address these gaps in the literature through more comprehensive benchmarks coupled with an analysis of why hierarchical structures in high-dimensional space may not add value. In the following section, we aim to reproduce the results of both Malkov and Yashunin (2016) and Lin and Zhao (2019) to confirm that we can independently replicate their findings with our own software implementation before proceeding to new experiments on larger-scale benchmark datasets.
3.2. Hubness in High Dimensional Spaces
Astute readers might observe that the HNSW graph construction algorithm described above does not explicitly enforce the small world property and instead adds edges between nodes based on their proximity in the metric space. The connection between proximity and the small world property arises due to hubness.
Specifically, hubness is a property of high-dimensional metric spaces where a small subset of points (the “hubs”) occur a disproportionate number of times in the near-neighbor lists of other points in the dataset (Radovanovic et al., 2010). In other words, a small fraction of nodes are highly connected to other points.
A related, but slightly different, notion of hubness that has been at the center of theoretical analysis is concentration of distances in high-dimensional spaces. It has been shown that, in expectation, the $\ell_{2}$ distance between independent and identically distributed (i.i.d) vectors grows with $\sqrt{d}$ while the variance tends to a constant as $d$ approaches infinity (Talagrand, 1994). As a result, the $\ell_{2}$ distance loses its discriminative power as $d$ increases. In fact, this concentration phenomenon is not only constrained to $\ell_{2}$ but also applies to other $\ell_{p}$ and even fractional norms (François et al., 2007).
Due to undesirable consequences of the hubness phenomenon, such as poor clustering quality, a large body of work has focused on hubness reduction strategies. For instance, (Zelnik-Manor and Perona, 2004) introduced local scaling which scales distances $d(\mathbb{x},\mathbb{y})$ by accounting for local neighborhood information. Interestingly, our work stands in contrast to this literature on hubness reduction by presenting a case study where hubs provide tangible value in an algorithmic setting, namely in accelerating greedy traversal in near neighbor proximity graphs. This result may be of independent interest to machine learning and algorithms researchers as well.
4. Reproduction of Prior Studies
@@
In this section, we present a replicability study using four flatnav NSW implementation. In particular, we revisit the experimental design of two prior works in the literature: the original 2016 HNSW paper of Malkov and Yashunin (2016) and a subsequent 2019 paper from Lin and Zhao (2019) that found limitations with the hierarchical component of HNSW. As we discussed in the previous section, these prior works possess limitations in experimental design, scope of benchmarking datasets, and a lack of analysis into understanding the results, which motivates our work in this paper. Nevertheless, we use these prior studies as a starting point to see if we can independently replicate these results via our own FlatNav implementation. Such a reproduction would both further validate the soundness of these previous experiments over the test of time as well as provide confirmation of the correctness of FlatNav before we proceed to new, larger-scale benchmarks.
@@
Following the same setups as Malkov and Yashunin (2016) and Lin and Zhao (2019) we generate a series of random vector datasets of varying dimensionality where each vector component is sampled uniformly at random from the range $[0,1)$. In particular, we consider dimensionalities of $d=4,8,16$ and 32. As in (Lin and Zhao, 2019), we set the number of near neighbors to retrieve to $k=1$ (departing from the default of $k=100$ we use elsewhere in this paper). We also tried including the sw-graph NSW baseline (Boytsov and Naidan, 2013) that Malkov and Yashunin (2016) used in their evaluation to benchmark against HNSW, but we were not able to run this older library successfully. However, we were able to replicate these prior findings using our own flatnav implementation which is conceptually identical to sw-graph but with more software optimizations to achieve engineering parity with hnswlib (detailed in Section 5.1).
@@
Figure 3. Median Latency vs. Recall of HNSW and FlatNav across dimensions $d=4,8,16,32$. We observe that the hierarchical structure accelerates search only when $d<32$, matching the findings of Lin and Zhao (2019). Our results demonstrating a significant advantage with HNSW on synthetic datasets with dimensionality $d=4$ and $d=8$ also match the findings of the original HNSW paper Malkov and Yashunin (2016)
@@
We also had the opportunity to discuss our findings with the lead author of (Malkov and Yashunin, 2016) who confirmed that the hierarchy provides a robust speedup on these low-dimensional datasets but noted the performance on higher dimensional vectors remained less clear, which further motivated us to take up the benchmarking study in the next section.
5. FlatNav Benchmarking Experiments
In this section, we report the results of our benchmarking study comparing the performance of flat HNSW search to hierarchical search on a suite of standard high-dimensional benchmark datasets drawn from real machine learning models.
A challenge in designing such benchmarking experiments is separating the confounding impact of performance engineering in the implementation of an approach from any purely algorithmic advantages of a method. We argue that this is a pitfall that has clouded several past benchmarking studies in ANN search. To circumvent this challenge, we fix the implementation in our experimental design such that the same code is used in constructing the base navigable small world graph. In particular, we use the popular hnswlib library as our baseline HNSW implementation. To benchmark the flat NSW index performance, we extract the bottom graph layer from hnswlib and reimplement the HNSW’s greedy search heuristic over this flat graph. To the best of our knowledge, we provide the first benchmarking study on the utility of HNSW’s hierarchy that makes an effort to achieve parity in the implementations of the two techniques to center the focus on the central algorithmic question.
5.1. Software Performance Optimizations
One major challenge in prior near neighbor benchmarking studies is the confounding effect of software performance optimizations when making conclusions about algorithmic efficacy. To address this issue, our implementation of a flat similarity search graph in flatnav closely follows the performance optimizations native to hnswlib. We use SIMD vector instructions to efficiently compute the distance between any two vectors, which is the key operation underpinning similarity search. Depending on the hardware type, we provide distance computations using SSE, AVX and AVX-512 intrinsics for float, uint8_t and int8_t data types. The choice of which intrinsics family to use is a compile-time choice, which is determined based on the dataset dimensionality and hardware support, similar to hnswlib. For all benchmark experiments, we use the float data type.
Furthermore, one common problem in constructing and querying graph-based indexes is the random data access pattern which can hamper latency and throughput. To improve search throughput, hnswlib implements prefetching using the _mm_prefetch intrinsic which reduces memory access latency by prefetching data into the L1 cache. Since caching can significantly affect latency metrics, we control for this performance optimization by leveraging the same intrinsic in the flatnav implementation so that our benchmarks strictly adhere to algorithmic differences.
5.2. Datasets and Compute
We utilize the benchmark datasets released through the popular leaderboards ANN Benchmarks (Aumüller et al., 2018) and Big ANN Benchmarks (Simhadri et al., 2022). The specific datasets and their associated statistics are presented in Table 1. For the Big ANN Benchmark datasets, we consider both the 10M and 100M collection of vectors for which the ground truth near neighbors have previously been computed and released. We did not experiment with the largest Big ANN datasets with 1 billion vectors since constructing HNSW indexes at this scale requires over 1.5TB of RAM, which exceeded our compute resources. Below, we include our benchmarking results for the four 100M-scale datasets available through Big ANN Benchmarks. We see that our flat HNSW implementation achieves performance parity with the hierarchical HNSW implementation.
@@ 5.2. Datasets and Compute
Table 1. Dataset Statistics. The datasets marked by $\dagger$ are from the BigANN benchmarks (Simhadri et al., 2022). The remaining are taken from ANN Benchmarks (Aumüller et al., 2018).
5.3. Latency Results
5.3.1. BigANN Benchmarks (Simhadri et al., 2022)
In Figures 4 and 5, we compare latency metrics for HNSW and FlatNav at the 50th and 99th percentile for the four 100M datasets from BigANN benchmarks listed in Table 1. All of our results support the conclusion that flatnav achieves nearly identical performance to hnswlib.
From the results in Figures 4 and 5, we observe that there is no consistent and discernable gap between FlatNav and HNSW in both the median and tail latency cases. These results suggest that the hierarchical structure of HNSW provides no tangible benefit on practical high-dimensional embedding datasets.
Figure 4. p50 Latency vs. Recall for HNSW and FlatNav over four 100M-scale benchmark datasets with dimensionality between 96 and 200. We observe that the flat NSW index achieves essentially identical performance to the hierarchical HNSW index.
Figure 5. p99 Latency vs. Recall for HNSW and FlatNav over four 100M-scale benchmark datasets with dimensionality between 96 and 200. As with the median case, we observe nearly identical performance between the two indexes.
5.3.2. ANN Benchmarks (Aumüller et al., 2018)
Moreover, we repeat the same experimental setup comparing HNSW and FlatNav on the ANN Benchmark datasets listed in Table 1. In Figures 6 and 7, we report the p50 and p99 latency of all of the non-GloVe ANN Benchmarks. Although these datasets are smaller in scale than the BigANN Benchmarks, we still see no discernible difference in latency between HNSW and FlatNav which supports our hypothesis that the vector dimensionality and not the size of the collection is the main driver of eliminating the need for hierarchical search in small world graphs. We see further evidence of this idea in Table 8, where we plot the median latency versus recall on the four GloVe benchmark datasets ranging in dimensionality from 25 to 200. Ultimately, these results, coupled with our reproducibility studies in the previous section, provide compelling evidence that for high-dimensional datasets, which are the standard in modern similarity search workloads, there is no apparent performance benefit to the hierarchical layers of HNSW.
Figure 6. p50 Latency vs. Recall for HNSW and FlatNav over five ANN Benchmark datasets ranging in dimensionality from 96 to 960. The flat NSW index again achieves practically indistinguishable performance from the hierarchical graph index.
Figure 7. p99 Latency vs. Recall for HNSW and FlatNav over five ANN Benchmark datasets ranging in dimensionality from 96 to 960.
Figure 8. p50 and p99 Latency vs. Recall for HNSW and FlatNav over the GloVe ANN Benchmark datasets with respective dimensionalities 25, 50, 100, and 200. We observe that HNSW has a minor advantage in low dimensions and that FlatNav has a minor advantage in high dimensions, but also that there is no clear difference in performance.
5.4. Memory
Dataset # Data hnswlib Memory flatnav Memory
BigANN 100M 183 113
Microsoft SpaceV 100M 104 85.5
Yandex DEEP 100M 100 60.7
Table 2. Peak Index Construction Memory in GBs. We observe that flatnav requires considerably less memory during construction compared to hnswlib.
In this section we measure the memory savings from removing the hierarchy by running a memory profiler during index construction. In terms of memory allocation, similarly to flatnav, hnswlib allocates static memory during index construction comprising base layer node allocation, a visited node list and a list of mutexes in the multi-threaded setting. Additionally, it also incurs memory cost attributable to the hierarchy, particularly maintaning the dynamically allocated links between nodes at each layer.
6. The Hub Highway Hypothesis
In the previous section, we established through a series of careful benchmarking experiments that there appears to be no discernible difference in performance between HNSW and its flat counterpart with no hierarchical graph structure. We now turn our attention to studying why the hierarchy appears to provide no benefit in the search process.
In our experiments, we observed that a small fraction of nodes appear in the set of near neighbors for a disproportionate number of other vectors. We thus conjecture that the hub structure prevalent in high-dimensional data performs the same functional role as the hierarchy.
Hypothesis 0 (Hub Highways).
In many applications, such as clustering, hubness is considered an undesirable property that harms algorithmic performance, motivating the need for hubness reduction techniques (Zelnik-Manor and Perona, 2004). On the contrary, our Hub Highway Hypothesis presents a case where hubness offers a distinct advantage in a practical algorithmic setting.
6.1. Intuition:
The formation of a hub highway can be understood through concentration of measure phenomena.
Concentration of distances: Under many data-generating distributions, the expected distances between randomly-drawn points converge to a constant value as dimension increases (Beyer et al., 1999). This suggests that in sufficiently high-dimensional spaces, we do not need to explicitly encourage the formation of “long-range” connections. Since all distances are approximately equal, all connections have similar length and a few edge traversals should span the graph of a high-dimensional dataset.
Concentration of measure: A well-known fact about high-dimensional geometry is that volume concentrates at the extrema of a shape (e.g., the surface of a ball or the faces of a hypercube). These portions of the distribution contain a disproportionate number of points and frequently contain the hubs of the metric space. For example, Low et. al. demonstrate that the the hubs of a uniform hypercube distribution are located at the corners (Low et al., 2013), which are exactly the points we would wish to connect via long-range connections in a hierarchical graph.
6.2. Methodology
  1. (1)
    Some nodes are visited by queries much more frequently than others. The relative popularity of these hub nodes is explained by the hubness phenomena that arises in high-dimensional datasets.
  2. (2)
    The hub nodes tend to be connected to each other, forming a well-connected subgraph of hubs.
  3. (3)
Empirical measures of hubness: To support the first part of our argument, we will need a formal definition and characterization of hubness.
Following (Radovanovic et al., 2010), let $\mathbb{x},\mathbb{x}_{1},\ldots,\mathbb{x}_{n}$ be a collection of vectors drawn from the same probability distribution with support $\mathcal{S}\subseteq\mathbb{R}^{d}$, and let $\phi:\mathcal{S}\times\mathcal{S}\to\mathbb{R}$ be a distance function.
Furthermore, for $1\leq i,k\leq n$, let $p_{i,k}$ be defined by
$p_{i,k}(\mathbb{x})=\begin{cases}1&\text{if $\mathbb{x}$ is among the $k$-NN set of $\mathbb{x}_{i}$ under $\phi$}\\ 0&\text{otherwise }\end{cases}$
Now let $N(\mathbb{x})$ be the random variable defined by
$N(\mathbb{x})\coloneqq\sum_{i=1}^{n}p_{i,k}(\mathbb{x})$
which represents the number of vectors that have $\mathbb{x}$ included in their $k$-nearest neighbors.
For any dataset $\mathcal{D}$, we can compute $N(\mathbb{x}),\mathbb{x}\in\mathcal{D}$, which yields a discrete distribution $N_{k}$. We are interested in the skewness of this distribution, given by
$S_{N_{k}}=\frac{\mathbb{E}\left[\left(N_{k}-\mu_{N_{k}}\right)^{3}\right]}{\sigma^{3}_{N_{k}}}$
This measure characterizes the asymmetry of the $k$-occurrence distribution $N_{k}$, and it is the metric most often used to estimate the presence of hubs. Thus, in our experiments we use $S_{N_{k}}$ to quantify the presence of hubness in a particular dataset. The more skewed the distribution of $N_{k}$, the more dataset $\mathcal{D}$ tends to have clusters of vectors (hubs) that keep regularly occurring in the $k$-NN sets of other vectors.
Table 3. Datasets
Dataset Dimensionality # Data # Queries
@@ 6.2. Methodology
IID Normal {16, 32, 64, 128, 256, 1024, 1536} 1M 10k
IID Normal {16, 32, 64, 128, 256, 1024, 1536} 1M 10k
Synthetic and ANN Benchmark Datasets: To illustrate evidence for the Hub Highway Hypothesis, we experiment with both real and synthetic datasets as shown in Table 3. In addition to a subset of ANN Benchmark datasets, we generate synthetic datasets by drawing vectors from the standard normal distribution. The choice of the normal distribution is important because, as (Low et al., 2013) observe, data sampled from the normal distribution exhibit a stronger hubness phenomenon due to strong density gradient than other distributions, such as sampling uniformly from a $d$-dimensional hypercube. Furthermore, since the hubness phenomenon is accentuated in high-dimensional spaces, we generate random data with increasing dimensionality which allows us to study the effect of these intrinsic highway structures that form as $d$ increases.
In the following three experiments we show empirical evidence supporting the Hub Highway Hypothesis. We start by examining the distribution of the number of times each node in the similarity search index is visited during search given a fixed number of queries, which we denote by $P_{m}(\mathbb{x}_{i})$ to indicate that node $\mathbb{x}_{i}$ is visited exactly $m$ times during a fixed number of $k$-NN queries for each one of the datasets in Table 3. We show that this distribution is skewed to the right, thus confirming that certain nodes (highway nodes) are visited a disproportionate number of times. Next, using this distribution, we selectively choose the most visited nodes in the similarity search graph to be the hub node clusters and show evidence that these nodes are more connected to each other than random nodes using hypothesis tests. In the last experiment, we show that not only are these highway nodes more connected, they also allow for faster graph traversals for queries, hence supporting the claim that we no longer need hierarchy in high dimensional vector search where fast query times evolve as a result of the presence of the hub-highways.
6.3. Skewness of the Node Access Distribution
As described above, for this study we consider the discrete distribution $P_{m}(\mathbb{x}_{i})$, $\mathbb{x}_{i}\in\mathcal{D}$, where $\mathcal{D}$ is one of the datasets in Table 3. For all these experiments we construct a HNSW index with parameters listed in Table 4 and import the base layer of the graph into FlatNav to construct a flat NSW index.
Table 4. Similarity search index parameters
$m$ $ef$-construction $ef$-search $k$
32 100 200 100
Figure 9. Log-normalized node access count distribution $P_{m}(\mathbb{x}_{i})$ for datasets using angular distance (left) and $\ell_{2}$ distance (right). Note that the x-axis is also log-normalized, so all histograms are skewed to the left – but the $\ell_{2}$ datasets have a much greater skew.
Figure 9 shows the log-normalized node access count distribution for different datasets. We observe that as the dimension $d$ increases, this distribution becomes more skewed to the right for $\ell_{2}$ distance-based datasets, indicating that a subset of nodes are visited much more often than the rest of the nodes in the graph. The cosine distance, on the other hand, is known to have anti-hub properties that prevent such a dramatic skew even for $d\in\{1024,1536\}$. This is consistent with the results observed by (Radovanovic et al., 2010). The increased skewness of the $P_{m}(\mathbb{x}_{i})$ distribution as $d$ increases suggests that highway nodes occur as a result of the intrinsic dimensionality of the datasets.

### v3

@@ A.1 Similarity Search
A.2 HNSW Overview
With this formalization of the ANNS problem, we will now briefly review the key elements of the HNSW algorithm, which is the central focus of our benchmarking study. As we alluded to previously, HNSW builds off of prior work in navigable small world graph indexes introduced in (Malkov et al., 2014). Small world graphs are a well-studied phenomenon in both computing and the social sciences and are primarily defined by the fact that the average length of a shortest path between two vertices is small (typically scaling logarithmically with the number of nodes in the network) (Travers & Milgram, 1977; Watts & Strogatz, 1998; Kleinberg, 2000). Small world graphs are also often characterized by the presence of well-connected hub nodes which we discuss further in the next section.
Specifically, the HNSW index is constructed in an iterative fashion. For a newly inserted element $x$, the algorithm will randomly select a maximum layer $l$ and then insert the new point into every layer up to $l$. This randomized process is executed with an exponentially decaying probability distribution such that, in expectation, each subsequent layer has exponentially more nodes than its predecessor. Within a layer, HNSW greedily adds edges between $x$ and its $M$ closest neighbors (where $M$ is a hyperparameter) where the neighbors consist of previously inserted points. This process then repeats in the subsequent layer below using the closest neighbors found in the prior graph as entry points. Through this process, the top layer of the hierarchy will be the coarsest directed graph, consisting of the fewest nodes and edges, and the bottom layer will be the densest and contain all of the nodes, each with connections to (up to) $M$ neighbors. As an additional, and important, optimization, HNSW also implements the pruning heuristic of (Arya & Mount, 1993) that will prune an edge from $u$ to $v$ if there exists another edge from $u$ to a neighbor $w$ of $v$ such that the distance from $u$ to $w$ is less than that of $u$ to $v$.
@@ A.2 HNSW Overview
32:     Return $M$ nodes based on selection heuristic in  (Arya & Mount, 1993)
@@
Appendix B Reproduction of Prior Studies
@@
In this section, we present a replicability study using four flatnav NSW implementation. In particular, we revisit the experimental design of two prior works in the literature: the original 2016 HNSW paper of (Malkov & Yashunin, 2016) and a subsequent 2019 paper from (Lin & Zhao, 2019) that found limitations with the hierarchical component of HNSW. As we discussed in the previous section, these prior works possess limitations in experimental design, scope of benchmarking datasets, and a lack of analysis into understanding the results, which motivates our work in this paper. Nevertheless, we use these prior studies as a starting point to see if we can independently replicate these results via our own FlatNav implementation. Such a reproduction would both further validate the soundness of these previous experiments over the test of time as well as provide confirmation of the correctness of FlatNav before we proceed to new, larger-scale benchmarks.
@@
Following the same setups as (Malkov & Yashunin, 2016) and (Lin & Zhao, 2019) we generate a series of random vector datasets of varying dimensionality where each vector component is sampled uniformly at random from the range $[0,1)$. In particular, we consider dimensionalities of $d=4,8,16$ and 32. As in (Lin & Zhao, 2019), we set the number of near neighbors to retrieve to $k=1$ (departing from the default of $k=100$ we use elsewhere in this paper). We also tried including the sw-graph NSW baseline (Boytsov & Naidan, 2013) that (Malkov & Yashunin, 2016) used in their evaluation to benchmark against HNSW, but we were not able to run this older library successfully. However, we were able to replicate these prior findings using our own flatnav implementation which is conceptually identical to sw-graph but with more software optimizations to achieve engineering parity with hnswlib.
@@
Figure 9: Median Latency vs. Recall of HNSW and FlatNav across dimensions $d=4,8,16,32$. We observe that the hierarchical structure accelerates search only when $d<32$, matching the findings of (Lin & Zhao, 2019). Our results demonstrating a significant advantage with HNSW on synthetic datasets with dimensionality $d=4$ and $d=8$ also match the findings of the original HNSW paper (Malkov & Yashunin, 2016)
@@
We also had the opportunity to discuss our findings with the lead author of (Malkov & Yashunin, 2016) who confirmed that the hierarchy provides a robust speedup on these low-dimensional datasets but noted the performance on higher dimensional vectors remained less clear, which further motivated us to take up the benchmarking study in the next section.
Appendix C Dataset Statistics
C.1 Benchmark Datasets for the Latency-Recall Tradeoff
Table 4 shows the different benchmark datasets used in Section 3.1.
@@ C.1 Benchmark Datasets for the Latency-Recall Tradeoff
Table 4: Dataset Statistics. The datasets marked by $\dagger$ are from the BigANN benchmarks (Simhadri et al., 2022). The remaining are taken from ANN Benchmarks (Aumüller et al., 2018).
C.2 Benchmark Datasets for the Hub-Highway Hypothesis Experiments
Table 5 details the various ANN and Big-ANN benchmark datasets as well as the synthetic datasets used in Section 4 for illustrating the empirical evidence of the hub-highway hypothesis.
Table 5: Hub-Highway Experimental Datasets
Dataset Dimensionality # Points # Queries
@@ C.2 Benchmark Datasets for the Hub-Highway Hypothesis Experiments
IID Normal {$2^{4}$, $2^{5}$, $2^{6}$, $2^{7}$, $2^{8}$, $2^{10}$, $1.5\cdot 2^{10}$} 1M 10k
IID Normal {$2^{4}$, $2^{5}$, $2^{6}$, $2^{7}$, $2^{8}$, $2^{10}$, $1.5\cdot 2^{10}$} 1M 10k
Appendix D Extended Discussion and Limitations
D.1 Extended Discussion
Small-World Graphs: The network science research community has known for decades that long-range connections and hubs induce the formation of “small-world” graphs that are easily traversed (Watts & Strogatz, 1998). This idea has been enormously influential in ANNS, providing the motivation for both NSW and HNSW (Malkov et al., 2012, 2014), but our results suggest that ANN graphs constructed over low-dimensional datasets may not in fact exhibit small-world properties. Because the $k$-occurrence distribution is near-uniform for intrinsically low-dimensional data distributions, a pure $k$NN graph (without pruning or long-range links) will not create hubs with a high in-degree. We believe that hierarchical structures are helpful in low dimensions because they help to induce hub behavior, by ensuring that search always begins from a small set of nodes. However, this is not necessary to produce hubs and induce small-world properties in high dimensions. The $k$NN graph construction process is sufficient on its own, because the hub highway emerges in high dimensions.
D.2 Limitations
Despite our notable empirical evidence for the hug-highway hypothesis and its significance in understanding graph-based ANN search, a principled understanding of this phenomenon from theoretical grounds is still lacking. Future research efforts can be directed towards explicit bounds on the probability of a query $q$ reaching high latency on a proximity graph $G=(V,E)$ consisting of hub nodes $H$.
There are several reasons why such a theoretical formulation is hard to attain. First, query latency is a metric that is the most perturbed by both engineering optimizations, such as SIMD operations, as well as the underlying hardware. Therefore, even if one attains a theoretical framework for bounding the expected query latency on a proximity graph, it is still very plausible that pure engineering optimizations could realize better performance in practice. Second, even if we control for engineering optimizations and the underlying hardware, it is not sufficient to consider the most obvious factors including graph complexity (i.e., $|V|$ and $|E|$) and the dimensionality of the underlying vector space $d$. Different graph construction procedures and pruning algorithms will induce different expected latency bounds. Future research directed to this theoretical understanding of the hub-highway hypothesis will prove to be invaluable to both theoreticians and practitioners alike.
Appendix E Extended Benchmarks
E.1 Building FlatNav from Scratch
In the main body of the paper, we present a series of results demonstrating that there is essentially no difference in performance between search over HNSW with a full hierarchy and search over the flat NSW base graph. As we discuss in Section 3, we fix the experimental design in the main body of the paper to extract the base NSW graph from the same hnswlib code that constructs the full HSNW index and use our flatnav implementation of the search algorithm to traverse the graph. In other words, we construct the full hierarchical graph with hnswlib and then extract the base layer as our separate flat graph index. We designed the benchmarking experiments in this manner to avoid any potential confounding effects from differences in code between the baseline HNSW graph construction and our own version.
However, this experimental setup raises a separate concern over whether the hierarchical component of HNSW might still be useful to construct the base graph index even if the hierarchy is not used during search. In this extended benchmark, we provide additional results to demonstrate that this is not the case. In particular, when we benchmark the full hierarchical HNSW index from hnswlib against a flat graph built from scratch with no hierarchy at all via our flatnav library, we observe identical results to what we report in Section 3. Thus, we can conclude that the hierarchical component of HNSW does not seem to be useful for high-dimensional workloads for either construction or search.
In Figures 11 and 11 we show the benchmarking results of HNSW and FlatNav with the latter built from scratch with no hierarchy even in construction. The results are identical to our findings in Figures 5 and 5 in Section 3 in the main body of the paper. For these extended benchmarks, we use a cloud server equipped with an AMD EPYC 9J14 96-Core Processor and 1 TB of RAM.
Figure 10: The p50 Latency vs Recall relationship between HNSW and FlatNav (constructed from scratch) is identical to the relationship between HNSW and FlatNav (base layer extracted from HNSW) shown in Figure 5.
Figure 11: The p50 Latency vs Recall relationship between HNSW and FlatNav (constructed from scratch) is identical to the relationship between HNSW and FlatNav (base layer extracted from HNSW) shown in Figure 5.
Appendix F Extended Hubness Experiments
F.1 Extending the Hub-Highway Hypothesis to LLM Embeddings
Figure 12: Log-normalized node access count distribution $P_{m}(\mathbb{x}_{i})$ for $\ell_{2}$ datasets along with MSMARCO embeddings.
To evaluate whether the Hub-Highway Hypothesis generalizes beyond sythetic gaussian-distributed and ANN benchmark datasets, we apply the same analysis to large language model (LLM) embeddings used in information retrieval. In particular, we examine the MSMARCO (Campos et al., 2016) dataset, a widely-used retrieval benchmark comprising millions of real-world queries and documents.
Data generation. We encode all training split queries in MSMARCO using the all-MiniLM-L6-v2 model from the SentenceTransformers (Reimers & Gurevych, 2019) library. This yields 384-dimensional vector representations. Using these embeddings, we construct a $k$-NN proximity graph (using $k=100$) and compute the node access distribution $P_{m}(\mathbb{x}_{i})$, defined as the number of times node $\mathbb{x}_{i}$ is visited across a fixed set of queries using HNSW beam search heuristic. As with all experiments, we fix the number of queries to be 10,000.
Results. Figure 12 shows the log-normalized node access frequency distribution for MSMARCO. The distribution exhibits a long-tail behavior similar to what we observed in synthetic $\ell_{2}$ datasets and the Big-ANN benchmarks, with a clear skew indicating that a small subset of nodes are accessed orders of magnitude more frequently than others. This supports our hypothesis that highway-like structures emerge naturally even in real-world retrieval embeddings generated by LLMs.
These findings suggest that the routing behavior of hub nodes is not an artifact of synthetic data or benchmark construction, but a general phenomenon in high-dimensional embedding spaces. This reinforces our claim that such hubs can effectively replace the hierarchy in modern ANN graph indexes.
F.2 No Hierarchy Memory Savings
Dataset Dataset Size hnswlib flatnav
BigANN 100M 183 113
Microsoft SpaceV 100M 104 85.5
Yandex DEEP 100M 100 60.7
Table 6: Peak Index Construction Memory in GBs. We observe that flatnav requires considerably less memory during construction compared to hnswlib.
Section 3 focused on demonstrating that we can remove the hierarchy in HNSW with impunity for latency benchmarks. Here we turn our attention to memory consumption and measure the memory savings from removing the hierarchy by running a memory profiler during index construction. In terms of memory allocation, similarly to flatnav, hnswlib allocates static memory during index construction comprising base layer node allocation, a visited node list and a list of mutexes in the multi-threaded setting. Additionally, it also incurs memory cost attributable to the hierarchy, particularly maintaning the dynamically allocated links between nodes at each layer.
F.3 Hub formation: Discerning Metric Space Hubness from Preferential Attachment